what is the amount of magnesium that would contain 1.20*10^24 particles? (Mg=24, Avogadro's constant+6.02*10^23)
1
Expert's answer
2015-08-11T09:50:45-0400
The number of moles of magnesium id defined by the equation:
µ = N/Na, where N- the number of particles and Na – Avogadro's constant;
µ = 1.20×10^24/ 6.02×10^23 mol^-1 = 1.9934 moles.
The mass of this amount of magnesium is:
m = µ × M, where M – the atomic weight of Mg
Thus, m = 1.9934 mol × 24 g/mol = 47.84 g